# SHARE: Shared Memory-Aware Open-Domain Long-Term Dialogue Dataset Constructed from Movie Script

### v1

@@ Abstract
Shared memories between two individuals strengthen their bond and are crucial for facilitating their ongoing conversations. This study aims to make long-term dialogue more engaging by leveraging these shared memories. To this end, we introduce a new long-term dialogue dataset named SHARE, constructed from movie scripts, which are a rich source of shared memories among various relationships. Our dialogue dataset contains the summaries of persona information and events of two individuals, as explicitly revealed in their conversation, along with implicitly extractable shared memories. We also introduce EPISODE, a long-term dialogue framework based on SHARE that utilizes shared experiences between individuals. Through experiments using SHARE, we demonstrate that shared memories between two individuals make long-term dialogues more engaging and sustainable, and that EPISODE effectively manages shared memories during dialogue. Our new dataset is publicly available at https://anonymous.4open.science/r/SHARE-AA1E/SHARE.json.
@@
Eunwon Kim         Chanho Park         Buru Chang Sogang University {eun1k, cksgh0984, buru}@sogang.ac.kr
@@ 1 Introduction
In this study, we introduce a new open-domain long-term dialogue dataset, SHARE, which includes not only personas and personal event information of individuals but also their shared memories. Collecting dialogue data that contain shared memory information presents the following challenges: First, similar to previous studies Xu et al. (2022b); Kwon et al. (2023), collecting dialogue data through crowdsourcing involves manually creating situations and then having crowd-workers role-play based on these situations, which incurs significant costs. Second, like other prior works Jang et al. (2023); Maharana et al. (2024), the approach of summarizing events explicitly evident in the dialogue between speakers using large language models (LLMs) misses their shared memories that can be implicitly inferred from dialogues. To address these challenges, we collect long-term dialogue data from movie scripts. In movie scripts, dialogues are used not only to describe characters, their relationships, and explicit events but also to convey shared memories that are not explicitly revealed in scenes Kozloff (2000). Thus, we consider movie scripts a rich source for collecting dialogue data that reflects shared memories between two individuals. Initially, we utilize a movie script parser Baruah and Narayanan (2023) to format the movie scripts into dialogues. Subsequently, we employ LLMs to summarize information about each speaker’s profile and short-term events, as well as to extract shared memories implicitly revealed in the dialogues. We then map this information to corresponding utterances for annotation.
@@ 3 Dataset Collection
Data source. We have collected a total of 1201 movie scripts from IMSDB [^1] https://imsdb.com/, DailyScript [^2] https://www.dailyscript.com/movie.html, and Simply Scripts. [^3] https://www.simplyscripts.com/ The collected scripts, encompassing a variety of genres including romance, comedy, and action, offer the advantage of presenting a diverse array of characters, relationships, and events.
@@
  • Persona information captures essential characteristics, including personality, occupation, and interests.
  • Personal event information covers transient details like impending deadlines or current health conditions.
  • Mutual event information captures significant interactions between the speakers, focusing on substantial events directly involving both individuals. Over time, these mutual events become new shared memories.
  • Shared memory refers to past memories that the two individuals have shared together prior to the current conversational context.
@@
Figure 2: An architecture of EPISODE. (1) Our framework selects memories from the memory set based on the current dialogue history to generate responses. (2) When the session ends, it extracts the information that needs to be stored and updates the existing memory set.
@@
Finally, the statistics of the new dataset are summarized in Table 2. We want to emphasize that 61.57% of episodes feature at least one shared memory in their conversations. This shows the significance of shared memory in maintaining meaningful dialogues between individuals who have known each other for a long time.
@@ 4.1 Task Definition
Let $\mathbf{e}_{(u,v)}$ = [$\mathbf{s}^{(1)},\mathbf{s}^{(2)},\cdots$] be an episode of two individuals $u$ and $v$ where each session $\mathbf{s}^{(i)}$ = [$\mathbf{u}_{1},\mathbf{v}_{1},\mathbf{u}_{2},\mathbf{v}_{2},\cdots,\mathbf{u}_{n},\mathbf{v}_{n}$] is a sequence of their utterances $\mathbf{u},\mathbf{v}$. The dialogue model $f(\cdot)$ generates a response $\tilde{\mathbf{v}}_{n}$ for the given context $\mathbf{c}$ = [$\mathbf{u}_{1},\mathbf{v}_{1},\cdots,\mathbf{u}_{n}$] and memory set $\mathcal{M}_{(u,v)}$ of the individuals as follows: $\tilde{\mathbf{v}}_{n}=f(\mathbf{c},\mathcal{M}_{(u,v)})$. In this paper, the memory set consists of persona information $\mathcal{P}_{u},\mathcal{P}_{v}$ and personal events $\mathcal{E}_{u},\mathcal{E}_{v}$ for each individual, and shared memories $\mathcal{S}_{(u,v)}$ between them, as $\mathcal{M}_{(u,v)}=\{\mathcal{P}_{u},\mathcal{P}_{v},\mathcal{E}_{u},\mathcal{E}_{v},\mathcal{S}_{(u,v)}\}$.
@@ 4.2 Response Generation
Memory selection. As long-term dialogue continues, the amount of persona information, personal events, and shared memories between the two individuals gradually increases. Using this expanding memory at each response generation stage is impractical. Therefore, we adopt a method that selects relevant memories for response generation based on the current context at each stage. This approach allows us to utilize memories from earlier conversations, supporting long-term dialogue. First, we design a prompt for memory selection and construct training data for the memory selector using this prompt and GPT-4o. To support dialogue situations that do not require specific memories, such as casual chitchat, we add ‘Everyday Language’ to the candidate pool where memories can be selected. The prompt instructs the model to select ‘Everyday Language’ when no suitable memory is available for the current context. Then, using Llama 3 Dubey et al. (2024), we train a memory selector to identify useful memories $\mathbf{m}$ from the current memory set $\mathcal{M}_{(u,v)}$ for response generation based on the given dialogue context $\mathbf{c}$.
Response generation. We use LLMs such as Gemma Team et al. (2024) and Llama Dubey et al. (2024) as the dialogue model $f(\cdot)$ to generate responses $\tilde{\mathbf{v}}_{n}$ appropriate to the given context $\mathbf{c}$ by utilizing the memories $\mathbf{m}$ selected in the previous step: $\tilde{\mathbf{v}}_{n}=f(\mathbf{c},\mathbf{m})$. To enable the dialogue model to incorporate the selected memories into response generation, we design specific prompts and collect training data based on these prompts to train the LLM. From this, the dialogue model generates personalized and consistent dialogues by providing responses that align with the memory while also considering the conversation history. The prompt used for dialogue generation is presented in Table 14, Table 15.
@@ 4.3 Memory Management
Information extraction. Once the current dialogue session ends, the information extraction process is initiated to extract distinct information from the session that may be memorable and useful for future dialogues, similar to how the SHARE dataset is constructed (see Section 3). The extracted information can be utilized in subsequent sessions. To reduce reliance on external APIs, we train Llama 3  Dubey et al. (2024) using the information extraction data collected during the construction of SHARE, and use it as the information extractor.
@@
LLaMA-3-Instruct (8B) LLaMA (zero-shot) 0.0122/0.0099 0.0997/0.0213 0.0923 0.8474 0.5458/0.8470 23.942
@@
LLaMA-3-Instruct (8B) SHARE (w/o shared memory) 2.5000 2.2438 1.5979 2.5375 2.2875 1.7021 2.5979 2.3538 1.6771
SHARE + ACCUMULATION 2.6125 2.2875 1.7354 1.0292 2.5833 2.2687 1.7167 1.2583 2.6583 2.3125 1.7271 1.3937
@@
Table 4: Experimental results of multi-session GPT evaluation.
@@ 5 Experiments
In this section, we conduct extensive experiments to valid the effectiveness of our new dataset and dialogue framework.
@@ 5.1 Experimental Setup
Backbone models. We employ state-of-the-art generative models, Gemma (2B) Team et al. (2024) and Llama-3-Instruct (8B) Dubey et al. (2024), as the backbone model of response generation.
@@
To verify that using shared memory makes long-term dialogues more engaging, we further conduct a GPT-4 assisted evaluation based on the following criteria:
  • Coherence(Coh): measuring the consistency and logical connection of responses within a session.
  • Engagingness(Eng): assessing the speaker’s ability to maintain the annotator’s interest for a long-term dialogue.
  • Closeness(Clo): determining how well the two participants know each other, considering the depth of their familiarity and understanding reflected in the dialogue.
  • Reflectiveness(Ref): evaluating how well the dialogue reflects the relationship indicated in the memory set.
Further details of these experiments are described in Appendix A.6
@@ 5.2 Experimental Results
  • SHARE w/o memory : This model is trained only with the dialogue from SHARE without using memories.
@@
  • SHARE w/ predicted memory : This model is trained with both the dialogue and annotated memories from SHARE. During inference, it generates the responses based on selected memories through the memory selection for the given dialogue context.
@@
  • SHARE w/ gold memory : This model uses the annotated gold memory to generate responses during inference.
@@
Table 3 shows the comparison of automatic evaluation metric results. Training with the SHARE dataset improves most performance metrics for generating the next response. In Llama 3, the BLEU-3 score increases by 0.0168 for SHARE w/o memory and 0.0267 for SHARE w/ predicted memory, while the ROUGE-1 score improves by 0.1146 and 0.1392, respectively, compared to the baseline performance of 0.0122. This reflects better performance in dialogue generation for using SHARE dataset. Beyond using the SHARE dataset, training the model with Shared Memory further enhances its performance. Specifically, SHARE w/ predicted memory achieves better perplexity, BLEU-3/4 and ROUGE-1/2 scores, while also demonstrating improved or comparable generation diversity in Dist-1/2. In summary, these results show that SHARE w/ predicted memory is a response generation model that effectively incorporates shared memory
Figure 3: Examples of dialogues with and without shared memory. The first and second utterances are the same in both dialogues, with the models generating responses from the third utterance onward. The conversation on the right builds upon Shared Memory from previous sessions. For instance, it includes discussions related to a driver’s license. In contrast, the conversation on the left lacks Shared Memory and revolves around general, everyday topics.
@@
Multi-Session Evaluation. As the number of sessions increases, the update model efficiently organizes the memory $\mathcal{M}_{(u,v)}$. To assess the performance of the updated model, we use a GPT-based evaluation approach. The test set represents as $\mathbf{e}_{(u,v)}=[\mathbf{s}^{(1)},\mathbf{s}^{(2)},\dots,\mathbf{s}^{(6)}]$, including more than six sessions, and a total of 96 episodes are selected. Each session $\mathbf{s}^{(i)}$ contains the utterances of two individuals, $u$ and $v$, structured as $[\mathbf{u}_{1},\mathbf{v}_{1},\mathbf{u}_{2},\mathbf{v}_{2},\dots,\mathbf{u}_{10},\mathbf{v}_{10}]$. The first two utterances are provided, and the model generates the following eight utterances, completing the session with a total of 10 sentences. Before generating the third session, the updated model integrates the information from the first and second sessions into $\mathcal{M}_{(u,v)}$, which is then used to generate the third session. After the session concludes, the relevant information is extracted. Starting from the fourth session, the model updates the memory with data from the previous two sessions to produce each new session, continuing this process until the sixth session.
@@
SHARE+EPISODE refers to our update model, which updates the memory throughout the sessions, while SHARE+ACCUMULATE accumulates all information as sessions progress without any updates. In the SHARE (w/o shared memory), $\mathcal{S}_{(u,v)}$ is excluded during memory selection. We evaluated the four metrics described in Section 5.1 on a scale from 0 to 3 points. Detailed prompts can be found in Table 17, Table 18, Table 19, Table 20.
Table 4 shows the evaluation of multi-session dialogues generated by the fine-tuned model. The high scores for each metric in every session indicate that the generated dialogues are engaging and coherent. Additionally, SHARE+EPISODE model consistently outperformed others in reflectiveness. Although the Llama 3 model showed lower scores in closeness in session 4, its performance improved in subsequent sessions, eventually surpassing the Accumulate model. This demonstrates the model’s effective use of shared experiences and its ability to clearly convey the relationship between the two participants. This success can be attributed to its strong conversational skills within the EPISODE framework and its efficient use of shared memory. We also conduct human evaluations under the same conditions. However, due to budget constraints, we were unable to expand the number of evaluators and the size of the test set sufficiently, so the results are included in the Appendix A.7.
Method Con Ref Eng
@@
Table 5: Results of different methods with and without shared memory.
EPISODE Evaluation. We evaluate the entire set of sessions together, rather than a single session. This is because, first, a good conversation should maintain consistency in the relationship between the two participants across all sessions. For example, if two characters are close in one session but suddenly become awkward in the next, it breaks the coherence of the dialogue. Second, a true long-term evaluation requires analyzing the dialogue from a broader perspective. If only one session is evaluated, important elements like references to past events that add interest could be missed. Thus, analyzing the entire conversation is necessary to accurately assess its coherence, Reflectiveness, and Engagingness.
Thus, we provide the full dialogue $\mathbf{e}_{(u,v)}=[\mathbf{s}^{(1)},\mathbf{s}^{(2)},\dots,\mathbf{s}^{(5)}]$, and evaluate the overall conversation in terms of Consistency, Reflectiveness, and Engagingness. In the test set, 131 episodes with more than 5 sessions were used for evaluation. Utterances were generated using the same method as in the multi-session evaluation. Here, consistency differs from coherence as it evaluates whether the relationship between the two speakers remains consistent throughout the entire episode. If any changes occur in the relationship, they must be properly supported by the dialogue. Detailed evaluation prompts are provided in the Appendix A.6.
Table 5 presents the evaluation of the full conversation sessions (1-5 sessions) generated by our EPISODE model. The results indicate that applying shared memory in each model leads to better performance compared to when it is not used. In terms of Reflectiveness, the SHARE+EPISODE model based on the Llama scores 2.3604 with shared memory, which is 0.0475 higher than the SHARE (w/o shared memory). This suggests that shared memories more effectively reveal the relationships between participants. For consistency, the Llama model with shared memory scores higher at 2.0729, showing a 0.0479 improvement over the SHARE model without shared memory. Overall, models with shared memory demonstrate stronger coherence in conversations compared to those without it. While the engagingness score for the Gemma model slightly decreases with shared memory, this is likely due to the inherent performance differences between the models. The Llama model initially performs better, and this advantage becomes more noticeable with shared memory. These results demonstrate that shared memories enhance both the reflectiveness and consistency of conversations, resulting in higher overall dialogue quality.

### v3

@@ Abstract
Shared memories between two individuals strengthen their bond and are crucial for facilitating their ongoing conversations. This study aims to make long-term dialogue more engaging by leveraging these shared memories. To this end, we introduce a new long-term dialogue dataset named SHARE, constructed from movie scripts, which are a rich source of shared memories among various relationships. Our dialogue dataset contains the summaries of persona information and events of two individuals, as explicitly revealed in their conversation, along with implicitly extractable shared memories. We also introduce EPISODE, a long-term dialogue framework based on SHARE that utilizes shared experiences between individuals. Through experiments using SHARE, we demonstrate that shared memories between two individuals make long-term dialogues more engaging and sustainable, and that EPISODE effectively manages shared memories during dialogue. Our dataset and code are available at https://github.com/e1kim/SHARE.
@@
Eunwon Kim1,thanks:   Equal contribution   Chanho Park1,11footnotemark: 1   Buru Chang2,thanks:   Corresponding author 1Sogang University   2Korea University {eun1k, cksgh0984}@sogang.ac.kr    buru_chang@korea.ac.kr
@@ 1 Introduction
In this study, we introduce a new open-domain long-term dialogue dataset, SHARE, which includes not only personas and personal event information but also information about memories shared between two speakers. Collecting dialogue data that contain shared memory information poses the following challenges: First, as in previous studies Xu et al. (2022b); Kwon et al. (2023), crowdsourced dialogue collection requires manually creating situations for crowd-workers to role-play, leading to high costs. Second, as in previous studies Jang et al. (2023); Maharana et al. (2024), attempts to extract events using large language models (LLMs) are primarily designed to summarize explicitly stated events within dialogues. Consequently, they often overlook implicitly inferred events that can be captured from the conversation context. To address these challenges, we compile long-term dialogue data from movie scripts. Movie scripts use dialogue to depict characters, their relationships, and explicit events, as well as to convey shared memories that are not directly revealed in scenes Kozloff (2000). Thus, movie scripts server as a valuable source for collecting dialogue data that reflects shared memories between two individuals. Initially, we use a movie script parser Baruah and Narayanan (2023) to format the movie scripts into dialogues. Subsequently, we utilize an LLM with a carefully designed prompt to summarize information about each speaker’s profile and short-term events, as well as to extract shared memories that are implicitly conveyed within the dialogues. We then map this information to the corresponding utterances for further annotation.
@@ 3 Dataset Collection
Data source. We have collected a total of 1201 movie scripts from IMSDB, [^1] https://imsdb.com/ DailyScript, [^2] https://www.dailyscript.com/movie.html and Simply Scripts. [^3] https://www.simplyscripts.com/ The collected scripts, encompassing a variety of genres including romance, comedy, and action, offer the advantage of presenting a diverse array of characters, relationships, and events.
@@
  • Persona information captures essential characteristics, including personality and interests.
  • Personal event information covers transient details like current health conditions.
  • Mutual event represents key events between two individuals that can be explicitly inferred from the current session, unfolding between them in real-time during the conversation. Over time, they evolve into shared memories.
  • Shared memory refers to past events that two individuals have previously shared before the current session. We capture these events that can be implicitly inferred from the dialogue.
@@
Figure 2: An architecture of EPISODE. Our framework selects memories from the memory set based on the current dialogue history to generate responses. When the session ends, it extracts the information that needs to be stored and updates the existing memory set.
@@
Finally, the statistics of the new dataset are summarized in Table 2. We want to emphasize that 61.57% of episodes feature at least one shared memory in their conversations. This shows the significance of shared memory in maintaining meaningful dialogues between individuals who have known each other for a long time. Interestingly, our dialogue dataset includes a diverse range of conversational styles, not only from everyday conversations but also from various genres, including fantasy. With dialogue models being increasingly applied beyond everyday conversations, such as in games Akoury et al. (2023) and films Chen et al. (2023), our dataset is a key resource for improving their adaptability across diverse dialogue styles.
@@ 4.1 Task Definition
Let $\mathbf{e}_{(u,v)}$ = [$\mathbf{s}^{(1)},\mathbf{s}^{(2)},\cdots$] be an episode of two individuals $u$ and $v$ where each session $\mathbf{s}^{(i)}$ = [$\mathbf{u}_{1},\mathbf{v}_{1},\mathbf{u}_{2},\mathbf{v}_{2},\cdots,\mathbf{u}_{n},\mathbf{v}_{n}$] is a sequence of their utterances $\mathbf{u},\mathbf{v}$. The dialogue model $f(\cdot)$ generates a response $\tilde{\mathbf{v}}_{n}$ for the given context $\mathbf{c}$ = [$\mathbf{u}_{1},\mathbf{v}_{1},\cdots,\mathbf{u}_{n}$] and memory set $\mathcal{M}_{(u,v)}$ of the individuals as follows: $\tilde{\mathbf{v}}_{n}=f(\mathbf{c},\mathcal{M}_{(u,v)})$. In this paper, the memory set consists of persona information $\mathcal{P}_{u},\mathcal{P}_{v}$ and personal events $\mathcal{E}_{u},\mathcal{E}_{v}$ for each individual, and shared memories $\mathcal{S}_{(u,v)}$ between them, as $\mathcal{M}_{(u,v)}=\{\mathcal{P}_{u},\mathcal{P}_{v},\mathcal{E}_{u},\mathcal{E}_{v},\mathcal{S}_{(u,v)}\}$. Note that all types of memory used in our study are extracted from conversations between two individuals. Therefore, we assume that both individuals have access to and can utilize these memories.
@@ 4.2 Response Generation
Memory selection. As long-term dialogue continues, the amount of persona information, personal events, and shared memories between the two individuals gradually increases. Using this expanding memory at each response generation stage is impractical. Therefore, we adopt a method that selects relevant memories for response generation based on the current context at each stage. This approach allows us to utilize memories from earlier conversations, supporting long-term dialogue. First, we design a prompt for memory selection and construct training data for the memory selection model. To support dialogue situations that do not require specific memories, such as casual chitchat, we add ‘Everyday Language’ to the candidate pool where memories can be selected. The prompt instructs the model to select ‘Everyday Language’ when no suitable memory is available for the current context. Then, using Llama-3-8B Dubey et al. (2024), we train a memory selector to identify useful memories $\mathbf{m}$ from the current memory set $\mathcal{M}_{(u,v)}$ for response generation based on the given context $\mathbf{c}$.
Response generation. We use Gemma Team et al. (2024) and Llama-3-8B Dubey et al. (2024) as the dialogue model $f(\cdot)$ to generate responses $\tilde{\mathbf{v}}_{n}$ appropriate to the given context $\mathbf{c}$ by utilizing the memories $\mathbf{m}$ selected in the previous step: $\tilde{\mathbf{v}}_{n}=f(\mathbf{c},\mathbf{m})$. To enable the dialogue model to incorporate the selected memories into response generation, we design specific prompts and collect training data based on these prompts to train the LLM. From this, the dialogue model generates personalized and consistent dialogues by providing responses that align with the memory while also considering the conversation history.
@@ 4.3 Memory Management
Information extraction. At the end of each dialogue session, an information extraction process is triggered to identify distinct and memorable details that may be useful for future interactions, following the approach used in the SHARE dataset. The extracted information is then incorporated into subsequent sessions. To minimize reliance on external APIs, we train Llama-3-8B on information extraction data collected during the construction of SHARE and utilize it as our information extractor.
@@
LLaMA-3.1-Instruct (8B) LLaMA (zero-shot) 0.0122/0.0099 0.0997/0.0213 0.0923 0.8474 0.5458/0.8470 23.942
@@
LLaMA-3.1-Instruct (8B) SHARE (w/o shared memory) 2.5000 2.2438 1.5979 2.5375 2.2875 1.7021 2.5979 2.3538 1.6771
SHARE + ACCUMULATION 2.6125 2.2875 1.7354 1.0292 2.5833 2.2687 1.7167 1.2583 2.5958 2.3125 1.7271 1.3937
SHARE + COMEDY 2.0104 1.2771 0.9854 2.0041 1.3000 0.9833 2.0625 1.2521 1.0042
SHARE + LLM-Rsum 2.5166 1.7646 1.2792 2.5709 1.7146 1.1583 2.6208 1.7063 1.2688
@@
Table 4: Experimental results of multi-session GPT-4o evaluation.
@@ 5.1 Experimental Setup
Backbone models. We employ state-of-the-art generative models, Gemma (2B) Team et al. (2024) and Llama-3.1-Instruct (8B) Dubey et al. (2024), as the backbone model of response generation.
@@
To verify that using shared memory makes long-term dialogues more engaging, we further conduct a GPT-4o assisted evaluation based on the following criteria: Coherence (Coh.): measures the consistency and logical connection of responses within a session. Engagingness (Eng.): assesses the speaker’s ability to maintain the annotator’s interest for a long-term dialogue. Closeness (Clo.): determines how well the two participants know each other, considering the depth of their familiarity and understanding. Reflectiveness (Ref.): evaluates how well the dialogue reflects the relationship indicated in the memory set.
Implementation details. We partition the SHARE dataset into an 8:1:1 ratio based on episodes, allocating the splits for training, validation, and testing in our long-term dialogue framework, EPISODE. Further details on these experiments are provided in Appendix A.6.
@@ 5.2 Experimental Results
  • SHARE w/o memory is trained only with the dialogue from SHARE without using memories.
@@
  • SHARE w/ predicted memory is trained with both the dialogue and annotated memories from SHARE. During inference, it generates the responses based on the memory selection for the given dialogue context.
  • SHARE w/ predicted individual memory is a baseline where each speacker independently manages their memory for response generation.
@@
  • SHARE w/ gold memory uses the gold memory to generate responses during inference.
@@
Table 3 shows the comparison of automatic evaluation metric results. Training with the SHARE dataset improves most performance metrics for generating the next response. In Llama-3, the BLEU-3/4 and ROUGE scores for SHARE w/ predicted memory responses are higher than those for SHARE w/ zero-shot and SHARE w/o memory This reflects better performance in dialogue generation for using the SHARE dataset, especially shared memory. Also, SHARE w/ predicted memory outperforms SHARE w/ predicted individual memory, demonstrating that incorporating the interlocutor’s persona and personal event information leads to more diverse and enriched conversations compared to their removal. In summary, these results show that SHARE w/ predicted memory is a response generation model that effectively incorporates shared memory.
@@
Multi-Session Evaluation. As the number of sessions increases, the update model efficiently organizes the memory $\mathcal{M}_{(u,v)}$. To assess the performance of the updated model, we use a GPT-based evaluation approach. The test set represents as $\mathbf{e}_{(u,v)}=[\mathbf{s}^{(1)},\mathbf{s}^{(2)},\dots,\mathbf{s}^{(6)}]$, including more than six sessions, and a total of 96 episodes are selected. Each session $\mathbf{s}^{(i)}$ contains the utterances of two individuals, $u$ and $v$, structured as $[\mathbf{u}_{1},\mathbf{v}_{1},\mathbf{u}_{2},\mathbf{v}_{2},\dots,\mathbf{u}_{10},\mathbf{v}_{10}]$. The first two utterances are provided, and the model generates the following eight utterances, completing the session with a total of 10 sentences. Before generating the third session, the updated model integrates the information from the first and second sessions into $\mathcal{M}_{(u,v)}$, which is then used to generate the third session. After the session concludes, the relevant information is extracted. From the fourth session, the model updates its memory with data from the past two sessions to generate each new session, continuing until the sixth.
@@
SHARE+EPISODE refers to our update model, which updates the memory throughout the sessions, while SHARE+ACCUMULATE accumulates all information as sessions progress without any updates. In the SHARE (w/o shared memory), $\mathcal{S}_{(u,v)}$ is excluded during memory selection. As a baseline, we use the COMEDY Chen et al. (2024) model and LLM-RsumWang et al. (2023). COMEDY analyzes previous session data to extract key details, such as significant events, user tendencies, and the nature of user-bot interactions, and then compresses them into a single memory. LLM-Rsum recursively summarizes and updates stored information at the end of each session. We evaluate the four metrics described in Section 5.1 on a scale from 0 to 3 points. More detailed information are in Appendix A.6.
Figure 3: Examples of dialogues with and without shared memory. The first and second utterances are the same in both dialogues, with the models generating responses from the third utterance onward. The conversation on the right builds upon shared memory from previous sessions.
Table 4 shows the evaluation of multi-session dialogues generated by the fine-tuned model. SHARE+EPISODE model consistently outperforms others in reflectiveness. Although the Llama-3 shows lower scores in closeness in session 4, its performance improves in subsequent sessions, eventually surpassing the Accumulate model. This highlights the model’s effective use of shared experiences and clear depiction of the participants’ relationship across multiple sessions.
Compared to the baselines, Gemma achieves higher metrics per session. Since LLM-Rsum recursively summarizes and updates stored information, evaluating the significance of individual memory units becomes challenging, which in turn leads to reduced Eng scores and lower Clo performance in conversations. The above results highlight the success of the EPISODE framework, which enhances conversational capabilities and effectively extracts both explicit and implicit information. Also, this approach proves to be more effective than previous approaches, which are primarily designed to summarize mainly explicitly stated events in dialogues, overlooking implicit experiences.
We also perform human evaluations under the same conditions, assessing key aspects such as coherence, reflectiveness, and engagingness of the generated dialogues. Detailed results of these evaluations are provided in Appendix A.7.
Method Con. Ref. Eng.
@@
Table 5: Comparison of different methods with and without shared memory based on GPT-4o evaluation.
EPISODE Evaluation. We evaluate the entire set of sessions together, rather than a single session. This is because, first, a good conversation should maintain consistency in the relationship between the two participants across all sessions. For example, if two characters are close in one session but suddenly become awkward in the next, it breaks the coherence of the dialogue. Second, a true long-term evaluation requires analyzing the dialogue from a broader perspective. If only one session is evaluated, important elements like references to past events that add interest could be missed. Thus, analyzing the entire conversation is necessary to accurately assess its Consistency (Con.), Reflectiveness (Ref.), and Engagingness (Eng.). Thus, we provide the full dialogue $\mathbf{e}_{(u,v)}=[\mathbf{s}^{(1)},\mathbf{s}^{(2)},\dots,\mathbf{s}^{(5)}]$, and evaluate the overall conversation. In the test set, 131 episodes with more than 5 sessions are used for evaluation. Utterances are generated using the same method as in the multi-session evaluation. Notably, Consistency differs from coherence in that it assesses whether the relationship between the two speakers remains stable and unchanged throughout the entire episode Detailed evaluation prompts are provided in the Appendix A.6.
Table 5 presents the evaluation of the full conversation sessions (1-5 sessions) generated by our EPISODE model. The results indicate that applying shared memory in each model leads to better performance compared to when it is not used. In Llama-3 and Gemma, SHARE+EPISODE achieves higher scores in Reflectiveness and Consistency compared to SHARE (w/o shared memory). This result indicates that shared memory enhances the model’s understanding of the relationship between the two participants, leading to more consistent dialogues. While the engagingness score for the Gemma slightly decreases with shared memory, this is likely due to the inherent performance differences between the models. The Llama-3 initially performs better, and this advantage becomes more pronounced with shared memory, further enhancing the overall quality of the conversation.